# “Sharing, Not Showing Off”: How BeReal Approaches Authentic Self-Presentation on Social Media Through Its Design

### v2

@@ 2.2. Authentic Self-Disclosure and Self-Presentation
Given that self-disclosure, especially in online contexts, involves extensive interpersonal interactions, it can be likened to a performance (Goffman, 2016) or an exhibition (Hogan, 2010) that symbolically communicate information about oneself to others (Blumer, 1986). As a result, self-presentation becomes an inherent part of this process. Self-disclosure is defined as “verbally communicating personal information about the self to another person” (Forgas, 2011) (p. 449), and reflects communicating a factual representation of oneself, regardless of its effect on one’s public image. In contrast, self-presentation is defined as “the goal-directed activity of controlling information to influence the impressions formed by an audience about the self” (Schlenker and Wowra, 2003) (p. 871). Thus, self-disclosure typically aligns with authenticity, whereas self-presentation may suggest a lack of authenticity. However, these two are not mutually exclusive (Johnson, 1981); revealing true aspects of oneself can still be compatible with self-presentation, as such revelations can also support a favorable public persona (Schlosser, 2020). We adopt the view in this paper that self-presentation is a strategic form of self-disclosure that predominates on social media, where users, acting as performers, anticipate positive outcomes from their disclosures.
@@ 6.1. The Social Media Self-Presentation Matrix (SMSM)
A 2x2 matrix with two types of self-presentation goals on the horizontal axis: “Self-Enhancing” (i.e., strategic social maneuvering goal) and “Self-Expressive” (i.e., accurate self-portrayal goal). On the vertical axis, we categorize the cognitive process of self-presentation as either “Automatic” (i.e., without conscious awareness, with minimal cognitive effort) or “Controlled” (i.e., deliberate, planned). The quadrant of Controlled and Self-Enhancing self-presentation is exemplified by the general strategy of highly manicured self-presentation on broadcasted social media platforms like Instagram. BeReal is intended to fall under the Automatic, Self-Expressive quadrant. We identify a design opportunity for social media platforms to move toward the Controlled, Self-Expressive quadrant. This quadrant would represent a more realistic approach toward authenticity, considering the intricacies of social media interactions, such as privacy issues, and the developmental characteristics of teens, such as heightened self-consciousness. We anticipate that a shift toward this quadrant could, in the long run, lead users or platform norms toward Automatic, Self-Expressive self-presentation, which is what BeReal aims to achieve.

### v3

@@ 2.2. Authentic Self-Disclosure and Self-Presentation
Given that self-disclosure, especially in online contexts, involves extensive interpersonal interactions, it can be likened to a performance (Goffman, 2016) or an exhibition (Hogan, 2010) that symbolically communicates information about oneself to others (Blumer, 1986). As a result, self-presentation becomes an inherent part of this process. Self-disclosure is defined as “verbally communicating personal information about the self to another person” (Forgas, 2011) (p. 449), and reflects communicating a factual representation of oneself, regardless of its effect on one’s public image. In contrast, self-presentation is defined as “the goal-directed activity of controlling information to influence the impressions formed by an audience about the self” (Schlenker and Wowra, 2003) (p. 871). Thus, self-disclosure typically aligns with authenticity, whereas self-presentation may suggest a lack of authenticity. However, these two are not mutually exclusive (Johnson, 1981); revealing true aspects of oneself can still be compatible with self-presentation, as such revelations can also support a favorable public persona (Schlosser, 2020). We adopt the view in this paper that self-presentation is a strategic form of self-disclosure that predominates on social media, where users, acting as performers, anticipate positive outcomes from their disclosures.
@@ 3. Background
[…]
Eight screenshots of BeReal, four at the top row and four at the bottom. Going in the order from the top left to the bottom right: The first image (a) shows a phone screen at 6:41 AM on Thursday, July 13, with a “Time to BeReal” notification at the bottom. The notification indicates that the user has 2 minutes left to capture a BeReal, with options to leave notifications on or turn them off. In the second image (b), the BeReal app is open, showing a live selfie of a woman smiling in an indoor space. A small inset image (e) in the top left corner captures the rear camera’s view of a computer screen. A timer at the top (b) reads “00:18.” The selfie (f) occupies most of the screen, with privacy options (“My friends only” and “Off”) and a “SEND” button visible at the bottom. The third image (c) shows a BeReal posted 14 hours late, with a timestamp of 1:05:01 PM. The BeReal includes an image of a computer screen with text, and the bottom of the image displays “6 retakes” (d) with the same privacy and sharing options as the previous image. In the fourth image (g), the BeReal shows a blank or blurry image with a red square in the upper left. A message in the center reads “Umm, anybody here?” The privacy settings and “SEND” button are at the bottom, similar to the other images. In the fifth image (h), the screen shows an “Instant RealMoji” feature with the prompt, “Slide and release to send.” A large circular preview displays a selfie of a woman giving a thumbs-up gesture. The instruction “Slide here to react” is written over the selfie, with a slider at the bottom prompting the user to send the reaction. The sixth image shows the user’s BeReal post, timestamped at 6:43:05 PM. The BeReal captures two images: a selfie of the user in the upper left corner and a view of their computer screen in the main frame, showing a document being worked on. At the bottom, a comment reads, “Hi BeReal paper! (paper emoji)(wave emoji),” with a “2 retakes” indicator visible. A small circular image of the user’s face (i) is positioned in the bottom left, next to the comment section. In the seventh image (j), the screen displays a blurred background with a message in the center: “Post to view. To view your friends’ BeReal, share yours with them.” Below this is a “Post a Late BeReal” button, prompting the user to submit their BeReal to unlock their friends’ posts. In the eight image (k), the user’s BeReal profile is shown. The profile displays a yellow icon with the letter “J” for the username “Jaewon.” Below the username, an “Add” button is visible with the instruction, “Add your real friends on BeReal.”
@@ 6.1. The Social Media Self-Presentation Matrix (SMSM)
The image is a diagram titled “The Social Media Self-Presentation Matrix (SMSM),” with two intersecting axes: (1) The vertical axis ranges from “Automatic” at the top to “Controlled” at the bottom. (2) The horizontal axis ranges from “Self-Enhancing / For Strategic Social Maneuvering” on the left to “Self-Expressive / For Accurate Self-Portrayal” on the right. Each quadrant represents a different type of self-presentation: 1. Top-left quadrant (Automatic + Self-Enhancing): Not labeled; 2. Top-right quadrant (Automatic + Self-Expressive): Represents self-presentation that is automatic and authentic, aimed at true self-expression. This is where “BeReal’s Intended Use Case” is located; 3. Bottom-left quadrant (Controlled + Self-Enhancing): Represents self-presentation that is controlled and strategic, typically involving careful curation for the purpose of social image enhancement. This is where “General Broadcasted Social Media” is positioned; 4. Bottom-right quadrant (Controlled + Self-Expressive): Represents self-presentation that is controlled and authentic, with an emphasis on accurate self-portrayal. This quadrant is labeled “(Untapped Design Opportunity)” in red, suggesting an unexplored area for social media design. Two dashed arrows from “General Broadcasted Social Media” to “(Untapped Design Opportunity),” and “(Untapped Design Opportunity)” to “BeReal’s Intended Use Case” indicate a two-step path.